OpenStudy (geometry_hater):

Evaluate this equation in exponential form Log base .0001 = -4

OpenStudy (anonymous):

Log base .0001 of what? =-4

OpenStudy (anonymous):

\[\log_{.0001} x=-4\]

OpenStudy (anonymous):

\[x=.0001^{-4}\]

OpenStudy (anonymous):

x=(10,000)^4=10^16

OpenStudy (anonymous):